OpenStudy (anonymous):

Can anybody help me find the Derivative of (4x)/(4x-6)? I have no idea how to do it other than the long formula.

OpenStudy (zzr0ck3r):

((4x)'(4x-6)-(4x-6)'(4x))/(4x-6)^2

OpenStudy (stamp):

\[d(\frac{4x}{4x-6})/dx=\]\[\frac{(4x-6)d(4x)-(4x)d(4x-6)}{(4x-6)(4x-6)}\]